# On face angles of tetrahedra with a given base

### v2

@@ 1. Introduction
$\Sigma(\triangle ABC):=\left\{\Bigl{(}\cos\overline{\alpha},\cos\overline{\beta},\cos\overline{\gamma}\Bigr{)}\in\mathbb{R}^{3}\,|\,ABCD\in\Omega(\triangle ABC)\right\}\,,$ (1)
where $\overline{\alpha}=\angle BDC$, $\overline{\beta}=\angle ADC$, and $\overline{\gamma}=\angle ADB$
@@
$\overline{\alpha}<\overline{\beta}+\overline{\gamma},\quad\overline{\beta}<\overline{\alpha}+\overline{\gamma},\quad\overline{\gamma}<\overline{\alpha}+\overline{\beta},\quad\overline{\alpha}+\overline{\beta}+\overline{\gamma}<2\pi$ (2)
@@
$\Upsilon(\triangle ABC):=\left\{\Bigl{(}x,y,z\Bigr{)}\in\mathbb{R}^{3}\,|\,ABCD\in\Omega(\triangle ABC)\right\}\,,$ (3)
where $x=d(D,C)$, $y=d(D,B)$, and $z=d(D,A)$
@@
$\mathbb{D}=\begin{vmatrix}0&x^{2}&y^{2}&z^{2}&1\\ x^{2}&0&\overline{z}^{2}&\overline{y}^{2}&1\\ y^{2}&\overline{z}^{2}&0&\overline{x}^{2}&1\\ z^{2}&\overline{y}^{2}&\overline{x}^{2}&0&1\\ 1&1&1&1&0\end{vmatrix}$ (4)
@@
$V^{2}\!\!\!$ =\!\!\!= $\frac{1}{36}x^{2}y^{2}z^{2}\Bigl{(}1+2\cos\overline{\alpha}\cdot\cos\overline{\beta}\cdot\cos\overline{\gamma}-\cos^{2}\overline{\alpha}-\cos^{2}\overline{\beta}-\cos^{2}\overline{\gamma}\Bigr{)}$ (5)
=\!\!\!= $\frac{1}{9}x^{2}y^{2}z^{2}\Bigl{(}\sin\frac{\overline{\alpha}+\overline{\beta}+\overline{\gamma}}{2}\cdot\sin\frac{\overline{\alpha}+\overline{\beta}-\overline{\gamma}}{2}\cdot\sin\frac{\overline{\alpha}+\overline{\gamma}-\overline{\beta}}{2}\cdot\sin\frac{\overline{\beta}+\overline{\gamma}-\overline{\alpha}}{2}\Bigr{)},$ (6)
@@
In our study, one subset of $\mathbb{E}^{3}$ plays an important role, namely, the right circular cylinder $\operatorname{Cyl}$ over the circle circumscribed around the base of $ABC$ (it suffices to consider only the part of this cylinder lying in $\Pi^{+}$).
@@
Section 5 is devoted to the study the image of $\Pi^{0}\setminus\{A,B,C\}$ under the map $F$. Note that $F(\Pi^{0}\setminus\{A,B,C\})\subset\mathbb{BP}$, the boundary of $\mathbb{P}$ from Theorem 1.
@@ 2. Using distance coordinates
$+(\overline{y}^{2}+\overline{z}^{2}-\overline{x}^{2})vw+(\overline{x}^{2}\overline{y}^{2}+\overline{x}^{2}\overline{z}^{2}-\overline{x}^{4})u+(\overline{x}^{2}\overline{y}^{2}+\overline{y}^{2}\overline{z}^{2}-\overline{y}^{4})v$ (7)
[…]
@@
It is easy to check that $(4,4)$-minor of $\mathbb{\overline{A}}$ (i.e. the determinant of the corresponding quadratic form) is zero. Therefore, the equality $\mathbb{D}=0$ defines some elliptic paraboloid, whereas the inequality $\mathbb{D}\geq 0$ described the convex hull of this elliptic paraboloid.
@@
[…]
@@
[…]
@@
$\overline{F}_{2}(x,y,z)$ $=$ $\frac{x^{2}+z^{2}-\overline{y}^{2}}{2xz}=\cos\angle CDA=\cos\overline{\beta},$ (8)
@@ 3. Cartesian coordinates and some general observations
Let us consider a Cartesian system in $\mathbb{R}^{3}$ such that $A=(a_{1},a_{2},0)$, $B=(b_{1},b_{2},0)$, $C=(c_{1},c_{2},0)$, and $D=(p,q,r)$. It is clear that the points $D=(p,q,r)$ and $D=(p,q,-r)$ give us two isometric tetrahedra with a common base $ABC$. Therefore, we may restrict our attention to the case $r\geq 0$.
@@
$\Pi^{+}=\{(p,q,r)\in\mathbb{R}^{3}\,|\,r>0\},\quad\Pi^{0}=\{(p,q,0)\in\mathbb{R}^{3}\},$ (10)
$\operatorname{GT}=\Pi^{+}\bigcup\bigl{(}\Pi^{0}\setminus\{A,B,C\}\bigr{)}.$ (11)
The latter set is the set of all tetrahedra with the base $ABC$ (we consider only the case when $D\not\in\{A,B,C\}$). We can also formally define the map $F:\operatorname{GT}\to\mathbb{R}^{3}$, using (3) and (2). Note, that
$F(p,q,r)=\bigl{(}\cos\overline{\alpha}=\cos\angle BDC,\cos\overline{\beta}=\cos\angle CDA,\cos\overline{\gamma}=\cos\angle ADB\bigr{)},$ (12)
@@
$\overrightarrow{\nabla}F_{2}=\overrightarrow{\nabla}\cos\overline{\beta}=\frac{x^{2}-z^{2}-\overline{y}^{2}}{2z^{3}x}\cdot\overrightarrow{DA}+\frac{z^{2}-x^{2}-\overline{y}^{2}}{2x^{3}z}\cdot\overrightarrow{DC},$ (13)
@@
$(\varphi,\psi)\mapsto\bigl{(}\cos(\varphi-\psi),\cos(\psi),\cos(\varphi)\bigr{)},\quad\varphi,\psi\in\mathbb{R}/2\pi\mathbb{Z}.$ (15)
@@
1) The surface $\mathbb{BP}$ is convex and it is smooth with positive Gaussian curvature in all points excepting the following four points: $(1,1,1)$, $(1,-1,-1)$, $(-1,1,-1)$, $(-1,-1,1)$.
2) $\mathbb{BP}$ is invariant under every permutation of the variables $x,y,z$ and under the map $(x,y,z)\mapsto(\pm x,\pm y,\pm z)$, where there are an even number of the signs ‘‘--’’ in the right side of the formula, hence, $\mathbb{BP}$ is invariant under the full tetrahedral group $S_{4}$.
3) The intersection of $\mathbb{BP}$ with the plane $z=1$ ($z=-1$) is the closed interval with the endpoints $(-1,-1,1)$ and $(1,1,1)$ (the closed interval with the endpoints $(-1,1,-1)$ and $(1,-1,-1)$, respectively).
4) The surface $\mathbb{BP}$ contains all edges of the tetrahedron with the vertices $(1,1,1)$, $(1,-1,-1)$, $(-1,1,-1)$, and $(-1,-1,1)$.
5) If $c=\operatorname{const}\in(-1,1)$, then the intersection of $\mathbb{BP}$ with the plane $z=c$ is an ellipse, which is invariant under the maps $(x,y)\mapsto(y,x)$ and $(x,y)\mapsto(-x,-y)$; in particular this ellipse is the unit circle for $c=0$.
Proof. The first property can be proved easily by the standard analytical methods. The second property is obvious. The third property follows from the equalities $(x-y)^{2}=0$ ($(x+y)^{2}=0$) for $z=1$ ($z=-1$, respectively). The next property follows easily from the previous three. Finally, it is clear that the equality $x^{2}-2cxy+y^{2}=1-c^{2}$ determines an ellipse with suitable symmetries.
@@
$\mathbb{T}_{KL}^{\,\alpha}=\left\{M\in{\mathbb{R}}^{3}\,|\,\angle KML=\alpha\right\}$ (16)
@@
$\frac{(p-k_{1})(p-l_{1})+(q-k_{2})(q-l_{2})+(r-k_{3})(r-l_{3})}{\sqrt{(p-k_{1})^{2}+(q-k_{2})^{2}+(r-k_{3})^{2}}\cdot\sqrt{(p-l_{1})^{2}+(q-l_{2})^{2}+(r-l_{3})^{2}}}=\cos(\alpha).$ (17)
Therefore, $\mathbb{T}_{KL}^{\,\alpha}$ is a subset of a suitable algebraic surface of degree $4$. If, say, $K=(-t,0,0)$ and $L=(t,0,0)$ for some $t>0$, then this surface is described by the equation
@@ 4. Degenerate points of $F$
Using Remark 1, we get that all $D$ from the plane $ABC$, $F$ is degenerate, hence, $D\in\operatorname{DF}$. Now we consider the case when $D$ is not in the plane $ABC$. In this case, the vectors $\overrightarrow{DA}$, $\overrightarrow{DB}$, and $\overrightarrow{DC}$ are linearly independent, hence $D\in\operatorname{DF}$ if and only if the determinant $|JF|$ of a $(3\times 3)$-matrix, which rows are the coordinates of the vectors $\overrightarrow{\nabla}F_{3}$, $\overrightarrow{\nabla}F_{2}$, $\overrightarrow{\nabla}F_{1}$ in the basis $\left(\overrightarrow{DC},\overrightarrow{DB},\overrightarrow{DA}\right)$, is 00.
@@
+.$+\Bigl{.}\overline{z}^{2}(z^{2}-y^{2})(x^{2}-z^{2})+\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}\Bigr{)}.$
@@
$\overline{x}^{2}(x^{2}-z^{2})(y^{2}-x^{2})+\overline{y}^{2}(z^{2}-y^{2})(y^{2}-x^{2})+\overline{z}^{2}(z^{2}-y^{2})(x^{2}-z^{2})+\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}=0,$ (20)
@@
$\frac{x^{2}-z^{2}}{\overline{y}^{2}}\cdot\frac{y^{2}-x^{2}}{\overline{z}^{2}}+\frac{z^{2}-y^{2}}{\overline{x}^{2}}\cdot\frac{y^{2}-x^{2}}{\overline{z}^{2}}+\frac{z^{2}-y^{2}}{\overline{x}^{2}}\cdot\frac{x^{2}-z^{2}}{\overline{y}^{2}}+1=0.$ (21)
@@
$\overline{x}^{2}+z^{2}-y^{2}=2\overline{x}\cdot z\cdot\cos\angle DAB,\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{y}^{2}+x^{2}-z^{2}=2\overline{y}\cdot x\cdot\cos\angle DCA,$
$\overline{z}^{2}+y^{2}-x^{2}=2\overline{z}\cdot y\cdot\cos\angle DBC,\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{x}^{2}+y^{2}-z^{2}=2\overline{x}\cdot y\cdot\cos\angle DBA,$
$\overline{y}^{2}+z^{2}-x^{2}=2\overline{y}\cdot z\cdot\cos\angle DAC,\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{z}^{2}+x^{2}-y^{2}=2\overline{z}\cdot x\cdot\cos\angle DCB,$
@@
$=-4xyz\overline{x}\leavevmode\nobreak\ \overline{y}\leavevmode\nobreak\ \overline{z}$
×$\times\Bigl{(}\cos\angle DAB\cdot\cos\angle DCA\cdot\cos\angle DBC+\cos\angle DBA\cdot\cos\angle DAC\cdot\cos\angle DCB\Bigr{)}=0.$
Let us recall some notation. Let $\overrightarrow{KL}$ and $\overrightarrow{MN}$ be collinear vectors. Denote by $\frac{\overline{KL}}{\overline{MN}}$ the quantity ±$\pm\frac{d(K,L)}{d(M,N)}$, where the plus sign is taken if the vectors $\overrightarrow{KL}$ and $\overrightarrow{MN}$ are directed in the same way, and the minus sign if the vectors are directed opposite to each other.
@@
$+(\overline{y}^{2}+\overline{z}^{2}-\overline{x}^{2})vw+\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}$ $=$ $0\,,$ (22)
@@
$\bigl{(}\overline{y}^{2}+\overline{z}^{2}-\overline{x}^{2}\bigr{)}\overline{x}^{2}\,u+\bigl{(}\overline{x}^{2}+\overline{z}^{2}-\overline{y}^{2}\bigr{)}\overline{y}^{2}\,v+\bigl{(}\overline{x}^{2}+\overline{y}^{2}-\overline{z}^{2}\bigr{)}\overline{z}^{2}\,w-2\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}=0.$ (23)

### v3

@@ 1. Introduction
(1) $\Sigma(\triangle ABC):=\left\{\Bigl(\cos\overline{\alpha},\cos\overline{\beta},\cos\overline{\gamma}\Bigr)\in\mathbb{R}^{3}\,|\,ABCD\in\Omega(\triangle ABC)\right\}\,$,
where $\overline{\alpha}=\angle BDC$, $\overline{\beta}=\angle ADC$, and $\overline{\gamma}=\angle ADB$ ?
@@
(2) $\overline{\alpha}<\overline{\beta}+\overline{\gamma},\quad\overline{\beta}<\overline{\alpha}+\overline{\gamma},\quad\overline{\gamma}<\overline{\alpha}+\overline{\beta},\quad\overline{\alpha}+\overline{\beta}+\overline{\gamma}<2\pi$
@@
(3) $\Upsilon(\triangle ABC):=\left\{\Bigl(x,y,z\Bigr)\in\mathbb{R}^{3}\,|\,ABCD\in\Omega(\triangle ABC)\right\}\,$,
where $x=d(D,C)$, $y=d(D,B)$, and $z=d(D,A)$ ?
@@
(4) $\mathbb{D}=\begin{vmatrix}0&x^{2}&y^{2}&z^{2}&1\\ x^{2}&0&\overline{z}^{2}&\overline{y}^{2}&1\\ y^{2}&\overline{z}^{2}&0&\overline{x}^{2}&1\\ z^{2}&\overline{y}^{2}&\overline{x}^{2}&0&1\\ 1&1&1&1&0\end{vmatrix}$
@@
(5) $V^{2}\!\!\!$ $=\!\!\!$ $\frac{1}{36}x^{2}y^{2}z^{2}\Bigl(1+2\cos\overline{\alpha}\cdot\cos\overline{\beta}\cdot\cos\overline{\gamma}-\cos^{2}\overline{\alpha}-\cos^{2}\overline{\beta}-\cos^{2}\overline{\gamma}\Bigr)$
(6) $=\!\!\!$ $\frac{1}{9}x^{2}y^{2}z^{2}\Bigl(\sin\frac{\overline{\alpha}+\overline{\beta}+\overline{\gamma}}{2}\cdot\sin\frac{\overline{\alpha}+\overline{\beta}-\overline{\gamma}}{2}\cdot\sin\frac{\overline{\alpha}+\overline{\gamma}-\overline{\beta}}{2}\cdot\sin\frac{\overline{\beta}+\overline{\gamma}-\overline{\alpha}}{2}\Bigr),$
@@
In our study, one subset of $\mathbb{E}^{3}$ plays an important role, namely, the right circular cylinder $\operatorname{Cyl}$ over the circle circumscribed around the base of ABCABC (it suffices to consider only the part of this cylinder lying in $\Pi^{+}$).
@@
Section 5 is devoted to the study of the image of $\Pi^{0}\setminus\{A,B,C\}$ under the map $F$. Note that $F(\Pi^{0}\setminus\{A,B,C\})\subset\mathbb{BP}$, the boundary of $\mathbb{P}$ from Theorem 1.
@@ 2. Using distance coordinates
(7) $+(\overline{y}^{2}+\overline{z}^{2}-\overline{x}^{2})vw+(\overline{x}^{2}\overline{y}^{2}+\overline{x}^{2}\overline{z}^{2}-\overline{x}^{4})u+(\overline{x}^{2}\overline{y}^{2}+\overline{y}^{2}\overline{z}^{2}-\overline{y}^{4})v$
$+(\overline{x}^{2}\overline{z}^{2}+\overline{y}^{2}\overline{z}^{2}-\overline{z}^{4})w-\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}>(\,\geq)\,0\,$,
@@
It is easy to check that $(4,4)$-minor of $\mathbb{\overline{A}}$ (i.e. the determinant of the corresponding quadratic form) is zero. Therefore, the equality $\mathbb{D}=0$ defines some elliptic paraboloid, whereas the inequality $\mathbb{D}\geq 0$ describes the convex hull of this elliptic paraboloid.
@@
$+(\overline{x}^{2}\overline{y}^{2}+\overline{x}^{2}\overline{z}^{2}-\overline{x}^{4})u+(\overline{x}^{2}\overline{y}^{2}+\overline{y}^{2}\overline{z}^{2}-\overline{y}^{4})v-\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}>(\,\geq)\,0$
@@
$\overline{x}^{2}\bigl(u-\overline{y}^{2}\bigr)^{2}-\bigl(\overline{x}^{2}+\overline{y}^{2}-\overline{z}^{2}\bigr)\bigl(u-\overline{y}^{2}\bigr)\bigl(v-\overline{x}^{2}\bigr)+\overline{y}^{2}\bigl(v-\overline{x}^{2}\bigr)^{2}<(\,\leq)\,0\,.$
@@
(8) $\overline{F}_{2}(x,y,z)$ $=$ $\frac{x^{2}+z^{2}-\overline{y}^{2}}{2xz}=\cos\angle CDA=\cos\overline{\beta},$
@@ 3. Cartesian coordinates and some general observations
Let us consider a Cartesian system in $\mathbb{R}^{3}$ such that $A=(a_{1},a_{2},0)$, $B=(b_{1},b_{2},0)$, $C=(c_{1},c_{2},0)$, and $D=(p,q,r)$. It is clear that the points $D=(p,q,r)$ and D=(p,q,r)D=(p,q,-r) give us two isometric tetrahedra with a common base $ABC$. Therefore, we may restrict our attention to the case $r\geq 0$.
@@
(10) $\Pi^{+}=\{(p,q,r)\in\mathbb{R}^{3}\,|\,r>0\},\quad\Pi^{0}=\{(p,q,0)\in\mathbb{R}^{3}\},$
(11) $\operatorname{GT}=\Pi^{+}\bigcup\bigl(\Pi^{0}\setminus\{A,B,C\}\bigr).$
The latter set is the set of all tetrahedra with the base ABCABC (we consider only the case when $D\not\in\{A,B,C\}$). We can also formally define the map $F:\operatorname{GT}\to\mathbb{R}^{3}$, using (3) and (2). Note, that
(12) $F(p,q,r)=\bigl(\cos\overline{\alpha}=\cos\angle BDC,\cos\overline{\beta}=\cos\angle CDA,\cos\overline{\gamma}=\cos\angle ADB\bigr),$
@@
(13) $\overrightarrow{\nabla}F_{2}=\overrightarrow{\nabla}\cos\overline{\beta}=\frac{x^{2}-z^{2}-\overline{y}^{2}}{2z^{3}x}\cdot\overrightarrow{DA}+\frac{z^{2}-x^{2}-\overline{y}^{2}}{2x^{3}z}\cdot\overrightarrow{DC},$
@@
(15) $(\varphi,\psi)\mapsto\bigl(\cos(\varphi-\psi),\cos(\psi),\cos(\varphi)\bigr),\quad\varphi,\psi\in\mathbb{R}/2\pi\mathbb{Z}.$
@@
1) $\mathbb{BP}$ is invariant under every permutation of the variables $x,y,z$ and under the map $(x,y,z)\mapsto(\pm x,\pm y,\pm z)$, where there are even number of the sign ‘‘-’’ in the right side of the formula.
2) The intersection of $\mathbb{BP}$ with the plane $z=1$ (z=1z=-1) is the closed interval with the endpoints (1,1,1)(-1,-1,1) and $(1,1,1)$ (the closet interval with the endpoints (1,1,1)(-1,1,-1) and (1,1,1)(1,-1,-1), respectively).
3) The surface $\mathbb{BP}$ contains all edges of the tetrahedron with the vertices $(1,1,1)$, (1,1,1)(1,-1,-1), (1,1,1)(-1,1,-1), and (1,1,1)(-1,-1,1).
4) If $c=\operatorname{const}\in(-1,1)$, then the intersection of $\mathbb{BP}$ with the plane $z=c$ is an ellipse, which is invariant under the maps $(x,y)\mapsto(y,x)$ and $(x,y)\mapsto(-x,-y)$; in particular this ellipse is the unit circle for $c=0$.
5) The surface $\mathbb{BP}$ is convex, it is smooth (with positive Gaussian curvature) at all points except the points $(1,1,1)$, (1,1,1)(1,-1,-1), (1,1,1)(-1,1,-1), (1,1,1)(-1,-1,1) (at all points except the edges of the tetrahedron with the vertices $(1,1,1)$, (1,1,1)(1,-1,-1), (1,1,1)(-1,1,-1), and (1,1,1)(-1,-1,1)).
Proof. The first property is obvious. The second property follows from the equalities $(x-y)^{2}=0$ ($(x+y)^{2}=0$) for $z=1$ (z=1z=-1, respectively). The third property is follows easily from the previous two. It is clear that the equality $x^{2}-2cxy+y^{2}=1-c^{2}$ determines an ellipse with suitable symmetries, which proves the fourth property. The last property can be proved easily by the standard analytical methods.
@@
(16) $\mathbb{T}_{KL}^{\,\alpha}=\left\{M\in{\mathbb{R}}^{3}\,|\,\angle KML=\alpha\right\}$
@@
(17) $\frac{(p-k_{1})(p-l_{1})+(q-k_{2})(q-l_{2})+(r-k_{3})(r-l_{3})}{\sqrt{(p-k_{1})^{2}+(q-k_{2})^{2}+(r-k_{3})^{2}}\cdot\sqrt{(p-l_{1})^{2}+(q-l_{2})^{2}+(r-l_{3})^{2}}}=\cos(\alpha).$
Therefore, $\mathbb{T}_{KL}^{\,\alpha}$ is a subset of a suitable algebraic surface of degree $4$. If, say, K=(t,0,0)K=(-t,0,0) and $L=(t,0,0)$ for some $t>0$, then this surface is described by the equation
@@ 4. Degenerate points of $F$
Using Remark 1, we get that all $D$ from the plane $ABC$, $F$ is degenerate, hence, $D\in\operatorname{DF}$. Now we consider the case when $D$ is not in the plane $ABC$. In this case, the vectors $\overrightarrow{DA}$, $\overrightarrow{DB}$, and $\overrightarrow{DC}$ are linearly independent, hence $D\in\operatorname{DF}$ if and only if the determinant $|JF|$ of a $(3\times 3)$-matrix, which rows are the coordinates of the vectors $\overrightarrow{\nabla}F_{3}$, $\overrightarrow{\nabla}F_{2}$, $\overrightarrow{\nabla}F_{1}$ in the basis $\left(\overrightarrow{DC},\overrightarrow{DB},\overrightarrow{DA}\right)$, is 0.
@@
$+\Bigl.\overline{z}^{2}(z^{2}-y^{2})(x^{2}-z^{2})+\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}\Bigr).$
@@
(20) $\overline{x}^{2}(x^{2}-z^{2})(y^{2}-x^{2})+\overline{y}^{2}(z^{2}-y^{2})(y^{2}-x^{2})+\overline{z}^{2}(z^{2}-y^{2})(x^{2}-z^{2})+\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}=0,$
@@
(21) $\frac{x^{2}-z^{2}}{\overline{y}^{2}}\cdot\frac{y^{2}-x^{2}}{\overline{z}^{2}}+\frac{z^{2}-y^{2}}{\overline{x}^{2}}\cdot\frac{y^{2}-x^{2}}{\overline{z}^{2}}+\frac{z^{2}-y^{2}}{\overline{x}^{2}}\cdot\frac{x^{2}-z^{2}}{\overline{y}^{2}}+1=0.$
@@
$\overline{x}^{2}+z^{2}-y^{2}=2\overline{x}\cdot z\cdot\cos\angle DAB,\penalty 10000\ \penalty 10000\ \overline{y}^{2}+x^{2}-z^{2}=2\overline{y}\cdot x\cdot\cos\angle DCA,$
$\overline{z}^{2}+y^{2}-x^{2}=2\overline{z}\cdot y\cdot\cos\angle DBC,\penalty 10000\ \penalty 10000\ \overline{x}^{2}+y^{2}-z^{2}=2\overline{x}\cdot y\cdot\cos\angle DBA,$
$\overline{y}^{2}+z^{2}-x^{2}=2\overline{y}\cdot z\cdot\cos\angle DAC,\penalty 10000\ \penalty 10000\ \overline{z}^{2}+x^{2}-y^{2}=2\overline{z}\cdot x\cdot\cos\angle DCB,$
@@
$=-4xyz\overline{x}\penalty 10000\ \overline{y}\penalty 10000\ \overline{z}$
$\times\Bigl(\cos\angle DAB\cdot\cos\angle DCA\cdot\cos\angle DBC+\cos\angle DBA\cdot\cos\angle DAC\cdot\cos\angle DCB\Bigr)=0.$
Let us recall some notation. Let $\overrightarrow{KL}$ and $\overrightarrow{MN}$ be collinear vectors. Denote by $\frac{\overline{KL}}{\overline{MN}}$ the quantity $\pm\frac{d(K,L)}{d(M,N)}$, where the plus sign is taken if the vectors $\overrightarrow{KL}$ and $\overrightarrow{MN}$ are directed in the same way, and the minus sign if the vectors are directed opposite to each other.
@@
(22) $+(\overline{y}^{2}+\overline{z}^{2}-\overline{x}^{2})vw+\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}$ $=$ $0\,$,
@@
(23) $\bigl(\overline{y}^{2}+\overline{z}^{2}-\overline{x}^{2}\bigr)\overline{x}^{2}\,u+\bigl(\overline{x}^{2}+\overline{z}^{2}-\overline{y}^{2}\bigr)\overline{y}^{2}\,v+\bigl(\overline{x}^{2}+\overline{y}^{2}-\overline{z}^{2}\bigr)\overline{z}^{2}\,w-2\overline{x}^{2}\overline{y}^{2}\overline{z}^{2}=0.$